OpenStudy (anonymous):

evaluate the expression log power 4 on bottom 1/1024

sam (.sam.):

\[\log_{4}\frac{1}{1024}\]

sam (.sam.):

\[\huge \log_{4}\frac{1}{4^5}\] \[=\huge \log_{4}4^{-5}\] =-5